# Weakly-Supervised Emotion Transition Learning for Diverse 3D Co-speech Gesture Generation

### v1

@@ Abstract
Generating vivid and emotional 3D co-speech gestures is crucial for virtual avatar animation in human-machine interaction applications. While the existing methods enable generating the gestures to follow a single emotion label, they overlook that long gesture sequence modeling with emotion transition is more practical in real scenes. In addition, the lack of large-scale available datasets with emotional transition speech and corresponding 3D human gestures also limits the addressing of this task. To fulfill this goal, we first incorporate the ChatGPT-4 and an audio inpainting approach to construct the high-fidelity emotion transition human speeches. Considering obtaining the realistic 3D pose annotations corresponding to the dynamically inpainted emotion transition audio is extremely difficult, we propose a novel weakly supervised training strategy to encourage authority gesture transitions. Specifically, to enhance the coordination of transition gestures w.r.t. different emotional ones, we model the temporal association representation between two different emotional gesture sequences as style guidance and infuse it into the transition generation. We further devise an emotion mixture mechanism that provides weak supervision based on a learnable mixed emotion label for transition gestures. Last, we present a keyframe sampler to supply effective initial posture cues in long sequences, enabling us to generate diverse gestures. Extensive experiments demonstrate that our method outperforms the state-of-the-art models constructed by adapting single emotion-conditioned counterparts on our newly defined emotion transition task and datasets.
@@ 1 Introduction
Co-speech gesture generation aims to synthesize vivid and emotional human postures coordinated with the audio input. These non-verbal behaviors serve as a key factor during human conversations that significantly facilitates the delivery of speech content [3, 12, 14]. Meanwhile, modeling co-speech gestures has a wide range of embodied AI applications in human-machine interaction [20, 25], robot assistants [7], and virtual/augmented reality (AR/VR) [9]. Conventionally, many researchers usually focus on synthesizing human upper body gestures consistent with speech audio [29, 49, 46, 44].
Nevertheless, except for a few recent works that generate the co-speech gestures of a single emotion category [26, 48, 2], previous works mostly focus on emotion-agnostic generation [46, 44]. Most of them overlook synthesizing the long sequence co-speech gestures with the emotion transitions, which are more practical in real-world scenes. For example, a person may not maintain a single emotion forever when communicating with others or in speech talking. In this work, we therefore introduce the task of speech-driven emotion transition for generating vivid and diverse 3D co-speech gestures, displayed in Figure 1. There are two main challenges in this task: 1) Datasets of 3D co-speech gestures synchronized with emotion transition speech audios are scarce. Creating such a dataset containing 3D human postures is difficult due to the lack of guidance from emotional experts and complex motion capture systems. 2) Modeling the plausible and temporal coherent co-speech gestures from one emotion to another in long sequences is difficult, especially in the transition duration.
@@
However, due to the dynamically generated transition transcripts, it is infeasible to construct the aligned realistic 3D pose annotation of human bodies. Hence, we solve the challenges of vivid co-speech gesture generation in a novel weakly supervised pattern, containing a motion transition infusion mechanism and an emotion mixture strategy. Specifically, in the motion transition infusion mechanism, we model the temporal correlation between the generated head and tail gesture features as style guidance representation. The style guidance representation provides motion transition cues that are infused into the transition embedding via an adaptive instance normalization (AdaIN) layer [16]. In this manner, we can effectively enhance the coordination of transition gestures w.r.t. two different emotional ones.
@@ 2.1 Co-speech Gesture Synthesis
Recently works focus on building the mapping directly from the input human speech and sequential gestures by exploiting the deep neural networks. They usually leverage multi-modality cues to facilitate the generation of co-speech gestures, associating with the speech audio [23, 49, 44], speaker identity [46], emotion [26, 2], and text transcript [29]. However, they overlook that emotion transition of the long sequential co-speech gesture modeling is much more practical in the real scenes. Moreover, since the lack of annotated 3D gestures corresponding to dynamically constructed transition speech, few of the above methods could be directly adapted to this new thought.
@@ 3.1 Emotion Transition Dataset Construction
Preliminary: Considering our key insight to modeling the co-speech gesture with emotion transition, we first split the existing aligned speech-gesture pairs [26, 33] into four-second clips. Then we randomly splice two clips from the same speaker to construct an emotion-transition candidate pair. The clip with neutral emotion is leveraged as head speech, and the other with various emotions is represented as tail speech. We leverage the dynamically synthesized two-second audio as a transition to combine the head and tail speeches. In the following, we briefly summarize our efforts in constructing the dynamic transition speech audio.
@@ 3.3 Weakly-supervised Emotion Transition
where indicates the transpose operation. With the help of the motion deformation representation, we obtain the two learnable emotion embeddings from the head and tail, respectively, as
@@ 4 Experiments
| FGD | FGD | BC | Diversity | FGD | FGD | BC | Diversity | |
| Seq2Seq [45] | 40.95 | 47.93 | 0.141 | 96.66 | 29.60 | 49.47 | 0.265 | 72.81 |
| S2G [11] | 25.56 | 37.04 | 0.671 | 98.26 | 18.16 | 41.63 | 0.824 | 76.82 |
| Trimodal [46] | 14.09 | 42.50 | 0.764 | 100.87 | 21.06 | 33.20 | 0.758 | 82.87 |
| CAMN [26] | 9.03 | 27.53 | 0.794 | 118.46 | 19.28 | 41.04 | 0.785 | 79.03 |
| HA2G [29] | 7.28 | 25.79 | 0.779 | 121.77 | 16.72 | 40.38 | 0.787 | 80.14 |
| DiffGesture [49] | 6.68 | 25.03 | 0.788 | 122.29 | 18.69 | 25.13 | 0.818 | 92.01 |
| Ours | 4.42 | 18.84 | 0.881 | 124.93 | 12.19 | 23.54 | 0.906 | 93.79 |
@@ 4.1 Datasets and Experimental Setting
| Baseline | MTIM | EMS | KFS | FGD | FGD | BC | Diversity | FGD | FGD | BC | Diversity |
| ✓ | 21.21 | 56.33 | 0.701 | 88.89 | 26.44 | 47.87 | 0.712 | 74.19 | |||
| ✓ | ✓ | 10.75 | 27.94 | 0.866 | 108.69 | 20.95 | 36.38 | 0.827 | 82.42 | ||
| ✓ | ✓ | ✓ | 5.69 | 21.70 | 0.878 | 112.80 | 14.02 | 26.21 | 0.900 | 87.21 | |
| ✓ | ✓ | ✓ | ✓ | 4.42 | 18.84 | 0.881 | 124.93 | 12.19 | 23.54 | 0.906 | 93.79 |
Evaluation Metrics: To fully evaluate the realism and diversity of the generated co-speech gestures, we introduce various metrics***More details are in the supplementary material.:
FGD: Fréchet Gesture Distance (FGD) [46] is utilized to measure the distribution distance between the realistic sequential gestures and generated ones. Since we only have the 3D joint annotation of the head/tail, we take the network architecture provided in [46, 29] to train the auto-encoder for distance computing on the two datasets, respectively. The FGD of transition gestures is calculated as the average value between the distribution distance of transition and head/tail, indicated as FGD. Similarly, FGD means the distance between the generated head/tail gestures and ground truth.
@@ 4.2 Quantitative Evaluation
As reported in Table 2, we adopt the FGD, FGD, BC, and Diversity for a well-rounded view of comparisons. Our method outperforms all the competitors by a large margin on both two datasets. Remarkably, on the TED-ETrans dataset, our method even achieves 34.8% (i.e., ) improvement over the sub-optimal counterparts in FGD. We observe both the DiffGesture [49] and ours synthesize the high-fidelity gestures of head/tail speech with much lower FGD than others. However, the DiffGesture shows worse performance on FGD due to lack of supervision. In terms of diversity, our simple yet effective keyframe sampler provides authority and diverse initial postures as the reference, thus enabling us to demonstrate diverse gesture styles compared to other counterparts. Moreover, we find that the diversity scores on BEAT-ETrans are much higher than those on TED-ETrans dataset. This can be attributed to the more complex human joints in the BEAT-ETrans dataset.
@@
Moreover, adopting the emotion mixture strategy ideally improves the performance of FGD on both two datasets. This indicates that the learnable emotion mixture wights can provide effective weak supervision by leveraging the pre-trained pose-based emotion classifier. Besides, we have observed significant improvement in the performance of FGD during this phase compared to the previous version. This aligns with our transformer backbone’s emphasis on modeling the sequential temporal correlations as a whole. The better transition gestures encourage our model to maintain better temporal consistency, thereby the head/tail gestures achieve better results.
@@ 4.3 Qualitative Evaluation
User Study: To further analyze the quality of results by various counterparts and ours, we conduct a user study by visualizing the results and inviting subjects. In particular, the subjects are required to rate the generated co-speech gestures from to (the higher, the better) in terms of naturalness, motion smoothness, and speech-gesture coherency. The results are demonstrated in Figure 4. Our method showcases the best performance compared with all the competitors. Especially in terms of motion smoothness, our method achieves noticeable advantages, indicating the effectiveness of our proposed motion transition infusion mechanism and the emotion mixture strategy.
@@ 9.1 Metric Calculation Details
Inspired by [29, 46], we take FGD to evaluate whether the generated gestures maintain realism with the ground truth ones in the perceptive of distribution. Conventionally, the feature extractor of FGD is calculated to embed overall sequential gestures into latent space and then utilize a decoder for reconstruction. However, since we do not have the ground truth of the transition gestures, we newly pre-train the feature extractor with the transition length . In the inference stage, FGD is calculated by averaging the distances between five randomly selected chunks of length from the head/tail and GT, respectively. Similarly, FGD is computed as the average value between the distance of transition and five randomly selected chunks of head/tail. We will release the code of our pipeline and evaluation metrics in the future.

### v3

@@ Abstract
Generating vivid and emotional 3D co-speech gestures is crucial for virtual avatar animation in human-machine interaction applications. While the existing methods enable generating the gestures to follow a single emotion label, they overlook that long gesture sequence modeling with emotion transition is more practical in real scenes. In addition, the lack of large-scale available datasets with emotional transition speech and corresponding 3D human gestures also limits the addressing of this task. To fulfill this goal, we first incorporate the ChatGPT-4 and an audio inpainting approach to construct the high-fidelity emotion transition human speeches. Considering obtaining the realistic 3D pose annotations corresponding to the dynamically inpainted emotion transition audio is extremely difficult, we propose a novel weakly supervised training strategy to encourage authority gesture transitions. Specifically, to enhance the coordination of transition gestures w.r.t. different emotional ones, we model the temporal association representation between two different emotional gesture sequences as style guidance and infuse it into the transition generation. We further devise an emotion mixture mechanism that provides weak supervision based on a learnable mixed emotion label for transition gestures. Last, we present a keyframe sampler to supply effective initial posture cues in long sequences, enabling us to generate diverse gestures. Extensive experiments demonstrate that our method outperforms the state-of-the-art models constructed by adapting single emotion-conditioned counterparts on our newly defined emotion transition task and datasets. Our code and dataset will be released on the project page: https://xingqunqi-lab.github.io/Emo-Transition-Gesture/.
@@ 1 Introduction
Co-speech gesture generation aims to synthesize vivid and emotional human postures coordinated with the audio input. These non-verbal behaviors serve as a key factor during human conversations that significantly facilitates the delivery of speech content [5, 14, 16]. Meanwhile, modeling co-speech gestures has a wide range of embodied AI applications in human-machine interaction [22, 27], robot assistants [9], and virtual/augmented reality (AR/VR) [11]. Conventionally, many researchers usually focus on synthesizing human upper body gestures consistent with speech audio [32, 53, 50, 48, 47].
Nevertheless, except for a few recent works that generate the co-speech gestures of a single emotion category [28, 52, 2, 3], previous works mostly focus on emotion-agnostic generation [50, 48, 4]. Most of them overlook synthesizing the long sequence co-speech gestures with the emotion transitions, which are more practical in real-world scenes. For example, a person may not maintain a single emotion forever when communicating with others or in speech talking. In this work, we therefore introduce the task of speech-driven emotion transition for generating vivid and diverse 3D co-speech gestures, displayed in Figure 1. There are two main challenges in this task: 1) Datasets of 3D co-speech gestures synchronized with emotion transition speech audios are scarce. Creating such a dataset containing 3D human postures is difficult due to the lack of guidance from emotional experts and complex motion capture systems. 2) Modeling the plausible and temporal coherent co-speech gestures from one emotion to another in long sequences is difficult, especially in the transition duration.
@@
However, due to the dynamically generated transition transcripts, it is infeasible to construct the aligned realistic 3D pose annotation of human bodies. Hence, we solve the challenges of vivid co-speech gesture generation in a novel weakly supervised pattern, containing a motion transition infusion mechanism and an emotion mixture strategy. Specifically, in the motion transition infusion mechanism, we model the temporal correlation between the generated head and tail gesture features as style guidance representation. The style guidance representation provides motion transition cues that are infused into the transition embedding via an adaptive instance normalization (AdaIN) layer [18]. Along with this manner, we can effectively enhance the coordination of transition gestures w.r.t. two different emotional ones.
@@ 2.1 Co-speech Gesture Synthesis
Recently works focus on building the mapping directly from the input human speech and sequential gestures by exploiting the deep neural networks. They usually leverage multi-modality cues to facilitate the generation of co-speech gestures, associating with the speech audio [25, 53, 48], speaker identity [50], emotion [28, 2, 31], and text transcript [32]. However, they overlook that emotion transition of the long sequential co-speech gesture modeling is much more practical in the real scenes. Moreover, since the lack of annotated 3D gestures corresponding to dynamically constructed transition speech, few of the above methods could be directly adapted to this new thought.
@@ 3.1 Emotion Transition Dataset Construction
Preliminary: Considering our key insight to modeling the co-speech gesture with emotion transition, we first split the existing aligned speech-gesture pairs [28, 36] into four-second clips. To diversify the datasets, we randomly splice two clips from the same speaker to construct an emotion-transition candidate pair. The clip with neutral emotion is leveraged as head speech, and the other with various emotions is represented as tail speech. We leverage the dynamically synthesized two-second audio as a transition to combine the head and tail speeches. In the following, we briefly summarize our efforts in constructing the dynamic transition speech audio.
@@ 3.3 Weakly-supervised Emotion Transition
where ′ indicates the transpose operation. With the help of the motion deformation representation, we obtain the two learnable emotion embeddings from the head and tail, respectively, as
@@ 4 Experiments
| FGDh+t | FGDtrans | BC | Diversity | FGDh+t | FGDtrans | BC | Diversity | |
| Seq2Seq [49] | 40.95 | 47.93 | 0.141 | 96.66±2.16 | 29.60 | 49.47 | 0.265 | 72.81±1.99 |
| S2G [13] | 25.56 | 37.04 | 0.671 | 98.26±2.04 | 18.16 | 41.63 | 0.824 | 76.82±2.32 |
| Trimodal [50] | 14.09 | 42.50 | 0.764 | 100.87±2.12 | 21.06 | 33.20 | 0.758 | 82.87±1.86 |
| CAMN [28] | 9.03 | 27.53 | 0.794 | 118.46±2.33 | 19.28 | 41.04 | 0.785 | 79.03±1.49 |
| HA2G [32] | 7.28 | 25.79 | 0.779 | 121.77±2.31 | 16.72 | 40.38 | 0.787 | 80.14±1.65 |
| DiffGesture [53] | 6.68 | 25.03 | 0.788 | 122.29±2.01 | 18.69 | 25.13 | 0.818 | 92.01±2.07 |
| Ours | 4.42 | 18.84 | 0.881 | 124.93±2.10 | 12.19 | 23.54 | 0.906 | 93.79±2.53 |
@@ 4.1 Datasets and Experimental Setting
| Baseline | MTIM | EMS | KFS | FGDh+t | FGDtrans | BC | Diversity | FGDh+t | FGDtrans | BC | Diversity |
| ✓ | 21.21 | 56.33 | 0.701 | 88.89±1.58 | 26.44 | 47.87 | 0.712 | 74.19±2.75 | |||
| ✓ | ✓ | 10.75 | 27.94 | 0.866 | 108.69±2.63 | 20.95 | 36.38 | 0.827 | 82.42±2.24 | ||
| ✓ | ✓ | ✓ | 5.69 | 21.70 | 0.878 | 112.80±1.95 | 14.02 | 26.21 | 0.900 | 87.21±2.12 | |
| ✓ | ✓ | ✓ | ✓ | 4.42 | 18.84 | 0.881 | 124.93±2.10 | 12.19 | 23.54 | 0.906 | 93.79±2.53 |
Evaluation Metrics: To fully evaluate the realism and diversity of the generated co-speech gestures, we introduce various metrics:
FGD: Fréchet Gesture Distance (FGD) [50] is utilized to measure the distribution distance between the realistic sequential gestures and generated ones. Since we only have the 3D joint annotation of the head/tail, we take the network architecture provided in [50, 32] to train the auto-encoder for distance computing on the two datasets, respectively. The FGD of transition gestures is calculated as the average value between the distribution distance of transition and head/tail, indicated as FGDtrans. Similarly, FGDh+t means the distance between the generated head/tail gestures and ground truth.
@@ 4.2 Quantitative Evaluation
As reported in Table 2, we adopt the FGDh+t, FGDtrans, BC, and Diversity for a well-rounded view of comparisons. Our method outperforms all the competitors by a large margin on both two datasets. Remarkably, on the TED-ETrans dataset, our method even achieves 34.8% (i.e., ) improvement over the sub-optimal counterparts in FGDh+t. We observe both the DiffGesture [53] and ours synthesize the high-fidelity gestures of head/tail speech with much lower FGDh+t than others. However, the DiffGesture shows worse performance on FGDtrans due to lack of supervision. In terms of diversity, our simple yet effective keyframe sampler provides authority and diverse initial postures as the reference, thus enabling us to demonstrate diverse gesture styles compared to other counterparts. Moreover, we find that the diversity scores on BEAT-ETrans are much higher than those on TED-ETrans dataset. This can be attributed to the more complex human joints in the BEAT-ETrans dataset.
@@
Moreover, adopting the emotion mixture strategy ideally improves the performance of FGDtrans on both two datasets. This indicates that the learnable emotion mixture wights can provide effective weak supervision by leveraging the pre-trained pose-based emotion classifier. Besides, we have observed significant improvement in the performance of FGDh+t during this phase compared to the previous version. This aligns with our transformer backbone’s emphasis on modeling the sequential temporal correlations as a whole. The better transition gestures encourage our model to maintain better temporal consistency, thereby the head/tail gestures achieve better results.
@@ 4.3 Qualitative Evaluation
User Study: To further analyze the quality of results by various counterparts and ours, we conduct a user study by visualizing the results and inviting subjects. For each emotion in both datasets, we randomly select 12 samples for each participant (7 emotions in BEAT-ETrans, 5 emotions in TED-ETrans). All the participants are recruited anonymously from schools including various majors. In particular, the subjects are required to rate the generated co-speech gestures from to (the higher, the better) in terms of naturalness, motion smoothness, and speech-gesture coherency. The results are demonstrated in Figure 4. Our method showcases the best performance compared with all the competitors. Especially in terms of motion smoothness, our method achieves noticeable advantages, indicating the effectiveness of our proposed motion transition infusion mechanism and the emotion mixture strategy.
@@ 5 Conclusion
Acknowledgements. This work is funded in part by the Theme-based Research Scheme grant (No.T45-205/21-N) and the InnoHK funding, Hong Kong SAR.
@@ 7 Dataset Construction
We first define GPT-4 prompt and evaluate sentence completeness confidence scores three times to select the best fit for semantic clarity. Moreover, we add a manual review on each transcript to drop the unnatural sentences, including 30 English native speakers’ evaluation of grammatical correctness/ logical coherence/ clarity of expression. We are unable to extract the transitional segments directly from lengthy videos due to the absence of speech recognizers for identifying multiple emotions within a single audio. Thus, we leverage the advanced LLM GPT-4 combined with the manual effort to construct the natural and smooth emotion transition datasets.
@@ 9.1 Metric Calculation Details
Inspired by [32, 50], we take FGD to evaluate whether the generated gestures maintain realism with the ground truth ones in the perceptive of distribution. Conventionally, the feature extractor of FGD is calculated to embed overall sequential gestures into latent space and then utilize a decoder for reconstruction. However, since we do not have the ground truth of the transition gestures, we newly pre-train the feature extractor with the transition length . In the inference stage, FGDh+t is calculated by averaging the distances between five randomly selected chunks of length from the head/tail and GT, respectively. Similarly, FGDtrans is computed as the average value between the distance of transition and five randomly selected chunks of head/tail. We will release the code of our pipeline and evaluation metrics in the future.
9.2 Additional Ablation Study Experiments
As reported in Table 4, after adding the adversarial loss, and BC achieve better results. This highly aligns with our motivation to ensure the temporal smoothness of the generated results. Inspired by BEAT, we leverage a pre-trained posture-based emotion classifier to evaluate the emotion transition effect in both datasets. As reported in Table 5, our method attains the best performance on emotion transition, which highly aligns with our visualization.
| Methods | BC | Diversity | ||
| Ours w/o | 15.31 | 32.72 | 0.802 | 79.64±4.58 |
| Ours | 12.19 | 23.54 | 0.906 | 93.79±2.53 |
| Models | BEAT-ETrans | TED-ETrans | ||||||
| FGDh+t | FGDtrans | BC | EmoACC | FGDh+t | FGDtrans | BC | EmoACC | |
| Seq2Seq | 40.95 | 47.93 | 0.141 | 57.50 | 29.60 | 49.47 | 0.265 | 56.20 |
| S2G | 25.56 | 37.04 | 0.671 | 60.69 | 18.16 | 41.63 | 0.824 | 58.88 |
| Trimodal | 14.09 | 42.50 | 0.764 | 69.81 | 21.06 | 33.20 | 0.758 | 63.82 |
| CAMN | 9.03 | 27.53 | 0.794 | 72.87 | 19.28 | 41.04 | 0.785 | 74.55 |
| HA2G | 7.28 | 25.79 | 0.779 | 73.98 | 16.72 | 40.38 | 0.787 | 80.74 |
| DiffGesture | 6.68 | 25.03 | 0.788 | 80.72 | 18.69 | 25.13 | 0.818 | 81.17 |
| Ours | 4.42 | 18.84 | 0.881 | 83.57 | 12.19 | 23.54 | 0.906 | 85.61 |